# LLMs as Planning Modelers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

### v1

@@ 3.1.2 Domain Modeling
Various works have executed domain modeling in a one-shot manner. Kalland (2024) combines a language and an automatic speech recognition model to generate PDDL specifications. To better understand cyber-attacks in real-time, CLLaMP (Oates et al. 2024) leverages LLMs to extract PDDL action models from Common Vulnerabilities and Exposures descriptions, finding that in-context (IC) examples are superior to CoT prompting. Zhang et al. (2024b) introduce PROC2PDDL, which proposes a Zone of Proximal Development prompt design—a variant of CoT (Vygotsky 1978). To mitigate misalignments with users requirements, Huang, Lipovetzky, and Cohn (2024) use multiple LLMs to generate candidate PDDL action schemas, which are then passed through a sentence encoder to compute the semantic relatedness of code and original NL descriptions.
@@
Guan et al. (2023) recognize the impracticality of LLMs generating fully functioning PDDL models in a single query (Kambhampati et al. 2024). Their framework LLM+DM outlines a generate-test-critique approach (Romera-Paredes et al. 2024; Trinh et al. 2024), leveraging multiple LLM calls to incrementally build key components of the domain by a dynamically generated predicate list. Similarly, Ishay and Lee (2025) introduce LLM+AL, which uses LLMs to generate action languages, but instead, in BC+ syntax (extension of Answer Set Programming). Shah (2024) presents LAMP, an extensive series of proposed algorithms that learn abstract PDDL domain models. Sinha (2024) proposes a structured prompt engineering approach to generate domain models in the Hierarchical Planning Definition Language (HPDL). Sikes et al. (2024b) explored decomposing NL information from operational models via Javascript functions into PDDL actions. BLADE (Liu et al. 2024a) bridges language-annotated human demonstrations and primitive action interfaces by tasking an LLM to define PDDL preconditions and effects conditioned on behaviors containing all possible sequences of contact primitive.
@@
In terms of closed-loop frameworks, ADA (Action Domain Acquisition) (Wong et al. 2023) tasks LLMs with generating candidate symbolic task decompositions, extracting undefined action names, and iteratively prompting for their definitions. COWP (Ding et al. 2023a) handles unforeseen situations in open-world planning by storing the robot’s closed-world state when planning fails, triggering a “Knowledge Acquirer” module that leverages LLMs to augment action preconditions and effects. Unlike COWP, which relies on predefined error factors, LASP (Chen et al. 2024) identifies potential errors from environmental observations, using an LLM to generate error causes in NL, suggesting action preconditions. Xie et al. (2024b) use fine-tuned LLMs for precondition and effect inference from NL actions, and semantic matching to validate actions by comparing inferred preconditions with the current world states.
@@
Oswald et al. (2024) shifts focus to assessing domain quality, addressing limitations of manual human evaluation (Hayton et al. 2020; Huang, Chen, and Zhang 2014). This study measures equivalence to the ground truth in terms of operational equivalence—whether reconstructed domains behave identically to the original by agreeing on the validity of action sequences as plans. To achieve this, the authors decompose ground truth PDDL actions into NL using an LLM, which then tasks them again to reconstruct PDDL domain models for quality assessment.
@@
Summary: Kambhampati et al. (2024); Wong et al. (2023) use incremental methods that iteratively generate and refine models, while incorporating real-world examples have shown to enhance contextual input accuracy (Oates et al. 2024; Ding et al. 2023a). The complexity of these frameworks demonstrate that constructing domains are inherently more challenging than task specification. However, current methods generate only a single domain, which risks misalignment with human expectations by failing to capture all necessary constraints. Future research should explore methods that generate multiple candidate domains (Huang, Lipovetzky, and Cohn 2024) and incorporate iterative refinement mechanisms to better align with user requirements.
@@ 3.1.3 Hybrid Modeling
Kelly et al. (2023) extract narrative planning domains and problems from input stories using a one-shot prompt, iterating with a second prompt conditioned on the planner’s error message until a successful plan is found. Smirnov et al. (2024) utilize pre-processing steps like JSON markup generation, consistency checks, and error correction loops. Their framework also includes a “reachability analysis” pipeline to extract feedback from flawed domains or unreachable problems, alongside a dependency analysis to check predicate usage across both files. ISR-LLM (Zhou et al. 2023) does not offer any feedback mechanisms to fix PDDL specifications; however, it does introduce self-refinement during the plan generation phase by incorporating the external validator tool, VAL (Howey, Long, and Fox 2004). Sakib and Sun (2024) generate multiple high-level task plans in Knowledge Graphs (KG), refine the network by removing unreliable components, and feed the task plan to an LLM to extract the PDDL domain and problem files for low-level robot skills. Conversely, DELTA (Liu et al. 2024b) initially generates PDDL files and a scene graph, followed by pruning unnecessary details from the graph to focus on relevant items.
@@
Huang and Zhang (2024) further support that LLMs are prone to one-shot generation errors, highlighting the need for intermediate representations before converting to PDDL. NL2Plan (Gestrin, Kuhlmann, and Seipp 2024) is the first domain-agnostic offline end-to-end NL planning system, requiring only minimal description and using pre-processing and automated common sense feedback to interface between the LLM and the user. LLM4CAP (da Silva et al. 2024) reduces manual effort, with an LLM-generated ontology being iteratively verified using an LLM to check for syntax errors, hallucinations, and missing elements. LLMFP (Hao, Zhang, and Fan 2024) translates goals, decision variables, and constraints into a JSON representation, which is then used to generate Python code for an SMT solver to produce plans without task-specific examples or external critics. NIPE (Ying et al. 2023) leverages LLMs as few-shot semantic parsers to generate conditional statements from spatial descriptions, guiding PDDL sampling and action model definition for Bayesian goal inference.
@@
Previous works tackle open-world settings, with MOPRHeus (Ye et al. 2024) focusing on human-in-the-loop long-horizon planning, introducing an anomaly detection mechanism to identify potential execution errors and update corresponding PDDL files to reflect changes in the world model. InterPret (Han et al. 2024) uses LLMs to enable robots to learn PDDL predicates and derive action schemas through interactive language feedback from non-expert users via Python perception APIs. Like other frameworks, Mahdavi et al. (2024) uses environmental interactions for evaluation and verification, starting with the LLM defining candidate PDDL problem files and domain sets, which are then refined through iterative cycles using their novel Exploration Walk (EW) method. Athalye et al. (2024) introduces pix2pred, a framework that leverages VLMs to propose predicates and determine their truth values in demonstrations. By bridging raw sensor data with symbolic planning, it structures the invented predicates to generalize across complex, long-horizon tasks, enabling the conversion of image data into planner-compatible representations, such as those based on PDDL.
@@
Instead of generating planning models to be used by external planners, AgentGen (Hu et al. 2024) uses LLMs to synthesize diverse environments and planning tasks (in languages such as PDDL) and their corresponding NL descriptions, targeted for LLM-based agent training. The authors found that LLMs instruction tuned with this dataset achieved significant improvement in both in-domain and out-of-domain planning tasks. As an exploration, (Hu et al. 2025) uses this dataset to run supervised fine-tuning with the LLaMA-3.1 models and found significant improvements in generating domain models–where larger models tend to benefit more from supervised fine-tuning.
@@
Summary: Complexities arise due to the need to coordinate between the domain and the problem. Human-in-the-loop interactions are frequently employed (Kelly et al. 2023), whereas other methods incorporate pre-processing steps–involving external tools like FastDownward and VAL (Zhou et al. 2023; Smirnov et al. 2024) or custom-designed rules (Mahdavi et al. 2024; Gestrin, Kuhlmann, and Seipp 2024). Yet, these linear pipelines risk cascading errors–such as the possibility of new objects in the task, prompting new PDDL types in the domain. Future work should focus on modularity, such as enabling dynamic integration of types and predicates in later stages of generation. This would result in more adaptable and error-tolerant planning systems.
@@ 3.2 Model Editing
Model edits have gained traction as a promising application functioning more as assistive tools than fully autonomous generative solutions. Insights into the likelihood of various model outcomes can support authors in refining models with greater efficiency toward an automated approach.
Gragera and Pozanco (2023) investigate the limitations of LLMs in repairing unsolvable tasks caused by incorrect task specifications, assessing the effectiveness of prompting in both PDDL and NL. As derivative work, Caglar et al. (2024) tackle the challenge of modifying model spaces beyond classical planning by exploring the effectiveness of LLMs in generating more plausible model edits, particularly concerning unsolvability and plan executability, either as a direct replacement or in augmentation of combinatorial search (CS) methods and manual authoring. Patil (2024) conduct a comprehensive study on using LLMs, with traditional error-checking methods, to detect and correct syntactic and semantic errors in PDDL domains, demonstrating that while LLMs excel at identifying and fixing syntactic issues, their performance in resolving semantic inconsistencies is less reliable. Sikes et al. (2024a) address the issues of planning models with domains containing a set of semantically equivalent state variables, which are syntactically different–provided with different labels. Specifically, they address failures in scenarios where information about the model is coming from heterogeneous sources.
Summary: Current research shows promise in leveraging LLMs to correct syntactic errors, but addressing semantic errors remains a significant challenge (Patil 2024). To overcome these issues, future work should explore post-hoc correction strategies. For instance, researchers could develop methodologies to analyze plan outputs after generation, identifying semantic gaps or inconsistencies through automated metrics or even human-in-the-loop evaluation systems. Such feedback can iteratively refine planning models, ensuring plans are both syntactically correct and semantically faithful.
@@ 3.3 Model Benchmarks
Non-deterministic output behaviors inherent in LLMs make it challenging to assess the quality of frameworks used in planning benchmarks. This heightens the importance of robustness, especially for evaluating LLMs’ ability to extract planning models (Behnke and Bercher 2024). However, this also introduces challenges of automating evaluation with these extracted models, determining levels of ambiguity in natural language inputs and assumptions, and monitoring fine-granular modes of failure in generation.
PlanBench (Valmeekam et al. 2023) aims to systematically evaluate LLM planning capabilities with an emphasis on cost-optimal planning and plan verification. ACPBench (Kokel et al. 2024) standardizes evaluation tasks and metrics for assessing reasoning about actions, changes (transitions), and planning—across 13 domains, on 22 SOTA language models including OpenAI’s o1 reasoning model. AutoPlanBench (Stein and Koller 2023) first converts PDDL planning benchmarks into NL via LLMs, and then tasks LLMs to produce a plan through various prompting techniques. Bohnet et al. (2024) provide a scalable benchmark suite in both PDDL and NL to measure the planning capabilities of LLMs in various strategies, as well as providing a mapping method for translating PDDL benchmarks to NL and measuring the performance of the generated benchmarks. Puerta-Merino et al. (2025) propose a road map and benchmark to address the gap of LLM integration in Hierarchical Planning (HP).
@@
To determine whether testing PDDL domains have been leaked to training data of LLMs, Mystery Blocksworld (Kambhampati et al. 2024) obfuscates the classic Blocksworld (Gupta, Efros, and Hebert 2010) planning problem by altering the named types to be syntactically nonsensical. ALFWorld and Household (Shridhar et al. 2021; Guan et al. 2023) tackles the complexities of real-world typical household environment that uses PDDL semantics to produce textual observations. TravelPlanner (Xie et al. 2024a) assesses language models’ abilities in planning through agent-based interactions in a travel-planning environment. Zheng et al. (2024) extend this work with Natural Plan, which evaluates LLMs on realistic planning and scheduling benchmarks using APIs.
@@
Previous LLM+AP benchmarks only focus on LLM plan generation capabilities, whereas Planetarium (Zuo et al. 2024) rigorously evaluates LLM-generated PDDL task specifications, highlighting two issues: (i) LLMs can produce valid code that misaligns with the original NL description, and (ii) evaluation sets often use NL descriptions too similar to the ground truth, reducing the task’s challenge. The benchmark assesses LLMs’ ability to generate PDDL problems across varying levels of abstraction and size. On the other hand, Text2World (Hu et al. 2025) introduces an automated pipeline for domain extraction and rigorous multi-criteria metrics that address the limitations of narrow domain scope (Liu et al. 2023a) and indirect evaluation methods in end-to-end plan assessments (Guan et al. 2023). Key metrics, including executability, structural similarity, and component-wise F1 scores, are employed while exploring state-of-the-art LLMs and fine-tuning techniques for generating and evaluating PDDL models.
@@
Summary: Assessing the quality of LLM-generated PDDL models (Zuo et al. 2024; Oswald et al. 2024) has made significant progress toward rigorous evaluation; however, the rapid leakage of training data to LLMs remains a major challenge, with (Hu et al. 2025) reporting high contamination rates in the evaluation domains from (Guan et al. 2023). “Mystifying” the domains (Kambhampati et al. 2024) does not address LLMs’ limitations in handling real-world modeling tasks. Other concerns arise that even with “mystified” domains, LLMs may still memorize code structures, rendering string obfuscation ineffective (Chen et al. 2025). Future work should explore solutions for routinely changing benchmark standards for domains, actively involving the planning community in its ongoing refinement. Khandelwal, Sheth, and Agostinelli (2024) proposes a tool for generating diverse and complex planning domains, which could serve as a foundation for such a benchmark.
4 (L2P) Library

### v2

@@ 3.1.2 Domain Modeling
Various works have executed domain modeling in a single query. To better understand cyber-attacks in real-time, CLLaMP (Oates et al., 2024) leverages LLMs to extract PDDL action models from Common Vulnerabilities and Exposures descriptions, finding that IC examples are superior to CoT prompting. Zhang et al. (2024b) introduce PROC2PDDL, which proposes a Zone of Proximal Development prompt design—a variant of CoT (Vygotsky, 1978). Sinha (2024) proposes a structured prompt engineering approach to generate domain models in the Hierarchical Planning Definition Language (HPDL). Huang et al. (2024b) use multiple LLM-generated candidate PDDL action schemas, which are then passed through a sentence encoder to compute the semantic relatedness of code and original NL descriptions. Following the candidate filtering approach, pix2pred (Athalye et al., 2024) leverages VLMs to propose predicates and determine their truth values in demonstrations.
@@
Guan et al. (2023) recognize the impracticality of LLMs generating fully functioning PDDL domains in a single call (Kambhampati et al., 2024). Their framework LLM+DM (Domain Model) outlines a generate-test-critique approach (Romera-Paredes et al., 2024; Trinh et al., 2024), leveraging multiple LLM calls to incrementally build key components of the domain by a dynamically generated predicate list. Similarly, Ishay and Lee (2025) introduce LLM+AL, which uses LLMs to generate action languages in BC+ syntax (extension of Answer Set Programming), while Sikes et al. (2024b) translates Javascript functions to PDDL in incremental stages. Shah (2024) presents LAMP, an extensive series of proposed algorithms that learn abstract PDDL domain models. BLADE (Liu et al., 2024a) bridges language-annotated human demonstrations and primitive action interfaces by tasking an LLM to define the PDDL action model preconditions and effects conditioned on behaviors containing all possible sequences of contact primitive and other behaviors preceding it.
@@
In terms of closed-loop frameworks, ADA (Action Domain Acquisition) (Wong et al., 2023) tasks LLMs with generating candidate symbolic task decompositions, extracting undefined action names, and iteratively prompting for their definitions. COWP (Ding et al., 2023a) handles unforeseen situations in open-world planning by storing the robot’s closed-world state when planning fails, triggering a “Knowledge Acquirer” module that leverages LLMs to augment action preconditions and effects. Unlike COWP, which relies on predefined error factors, LASP (Chen et al., 2024) identifies potential errors from environmental observations, using an LLM to generate error causes in NL, suggesting action preconditions. Xie et al. (2024b) use fine-tuned LLMs for precondition and effect inference from NL actions, and semantic matching to validate actions by comparing inferred preconditions with the current world states.
@@
To evaluate domain quality, Oswald et al. (2024) addresses limitations of manual human evaluation (Hayton et al., 2020; Huang et al., 2014) and string-based comparison methods that assess similarity to ground truth. This study measures equivalence to the ground truth in terms of operational equivalence—whether reconstructed domains behave identically to the original by agreeing on the validity of action sequences as plans. To achieve this, the authors decompose ground truth PDDL actions into NL using an LLM, which then tasks them again to reconstruct PDDL domain models for quality assessment.
@@
Summary and Future Directions: Kambhampati et al. (2024); Wong et al. (2023) use incremental methods that iteratively refine models. Real-world examples have shown to enhance contextual output accuracy (Oates et al., 2024; Ding et al., 2023a). The complexity of these frameworks demonstrate that constructing domains are inherently more challenging than task specification. However, by generating and relying on a single domain model, current methods risk rendering the entire planning process invalid if that model fails to capture implicit user constraints. Future approaches should consider generating multiple candidate domains—or specific components, such as predicate definitions—to better accommodate ambiguity and uncertainty in user intent (Huang et al., 2024b; Athalye et al., 2024).
@@ 3.1.3 Hybrid Modeling
Hybrid modeling combines PDDL domain and problem systems. Kelly et al. (2023) extract narrative planning domains and problems from input stories using a one-shot prompt, iterating with a second prompt conditioned on the planner’s error message until a successful plan is found. ISR-LLM (Zhou et al., 2023) does not offer any feedback mechanisms to fix PDDL specifications; however, it does introduce self-refinement during the plan generation phase by incorporating the external validator tool, VAL (Howey et al., 2004). Sakib and Sun (2024) generate multiple high-level task plans in Knowledge Graphs (KG), prunes unnecessary components, and feed the task plan to an LLM to extract the PDDL domain and problem files for low-level robot skills. Conversely, DELTA (Liu et al., 2024b) initially generates PDDL files and a scene graph, followed by pruning unnecessary details from the graph to focus on relevant items.
@@
Huang and Zhang (2024) further support that LLMs are prone to one-shot generation errors, highlighting the need for intermediate representations before converting to PDDL. NL2Plan (Gestrin et al., 2024) is the first domain-agnostic offline end-to-end NL planning system, requiring only minimal description and using pre-processing and automated common sense feedback to interface between the LLM and the user. Smirnov et al. (2024) utilize pre-processing steps like JSON markup generation, consistency checks, and error correction loops. Their framework also includes a “reachability analysis” pipeline to extract feedback from flawed domains or unreachable problems, alongside a dependency analysis to check predicate usage across both files. LLM4CAP (da Silva et al., 2024) reduces manual effort, with an LLM-generated ontology being iteratively verified using an LLM to check for syntax errors, hallucinations, and missing elements. LLMFP (Hao et al., 2024) translates goals, decision variables, and constraints into a JSON representation, which is then used to generate Python code for an SMT solver to produce plans without task-specific examples or external critics. NIPE (Ying et al., 2023) leverages LLMs as few-shot semantic parsers to generate conditional statements from spatial descriptions, guiding PDDL sampling and action model definition for Bayesian goal inference.
@@
For real-world grounding, MORPHeus (Ye et al., 2024) focuses on human-in-the-loop long-horizon planning, introducing an anomaly detection mechanism to identify potential execution errors and update corresponding PDDL files to reflect changes in the world model. InterPret (Han et al., 2024) uses LLMs to enable robots to learn PDDL predicates and derive action schemas through interactive language feedback from non-expert users via Python perception APIs. Mahdavi et al. (2024) uses environmental interactions for evaluation and verification, starting with the LLM defining candidate PDDL problem files and domain sets, which are then refined through iterative cycles using their novel Exploration Walk (EW) method.
@@
Instead of generating models for external planners, AgentGen (Hu et al., 2024) uses LLMs to synthesize diverse PDDL tasks and NL descriptions for training LLM-based agents. Their work demonstrate that instruction-tuned models show significant gains in both in-domain and out-of-domain planning tasks. Hu et al. (2025) further fine-tune LLaMA-3.1 on this dataset, finding notable improvements in domain model generation, especially for larger models.
@@
Summary and Future Directions: Complexities arise when coordinating the domain and respective problem. Human-in-the-loop interactions are frequently employed (Kelly et al., 2023), whereas other methods incorporate pre-processing steps—involving external tools like FastDownward and VAL (Zhou et al., 2023; Smirnov et al., 2024) or custom-designed rules (Mahdavi et al., 2024; Gestrin et al., 2024). These linear pipelines risk cascading errors, such as the possibility of new objects in the later stages of the task, prompting new PDDL types in the domain. Future work should focus on modularity, such as enabling dynamic integration of types and predicates in later stages of generation. This would result in more adaptable and error-tolerant planning systems.
@@ 3.2 Model Editing
The use of LLMs serving more as assistive tools than fully autonomous generative solutions has shown promising applications for LLM+AP integration. Understanding LLM-editing decisions in refining specifications can support authors with greater efficiency toward an automated approach.
Gragera and Pozanco (2023) investigate the limitations of LLMs in repairing unsolvable tasks caused by incorrect task specifications, assessing the effectiveness of prompting in both PDDL and NL. Patil (2024) conduct a comprehensive study on using LLMs, with traditional error-checking methods, to detect and correct syntactic and semantic errors in PDDL domains, demonstrating that LLMs excel at syntax correction but are less reliable with semantic inconsistencies. Sikes et al. (2024a) addresses planning model failures caused by semantically equivalent but syntactically distinct state variables, a common issue when integrating information from heterogeneous sources. Their approach introduces meta-actions to bridge these mismatches and iteratively refines the model to ensure valid plan generation. Caglar et al. (2024) address the challenge of modifying model spaces beyond classical planning by evaluating how effectively LLMs generate plausible model edits—especially to fix unsolvability and plan executability—to support combinatorial search and manual methods.
Summary and Future Directions: Current research shows promise in using LLMs to correct syntactic errors, but addressing semantic errors remains a significant challenge (Patil, 2024) leading to non-executable or semantically inconsistent plans. Future work should explore post-hoc correction strategies. For instance, researchers could explore strategies to analyze plan outputs, identifying semantic inconsistencies through automated metrics or human evaluation systems as grounded feedback for error-correction.
@@ 3.3 Model Benchmarks
LLMs, with their non-deterministic output behaviors, make it challenging to assess the quality of frameworks used in planning benchmarks. This heightens the importance of robustness, especially for evaluating LLMs’ ability to extract planning models (Behnke and Bercher, 2024).
LLM+AP benchmarks typically fall into two categories: (1) evaluating LLMs’ direct planning abilities; (2) assessing the quality of planning specifications produced by LLMs. While this survey focuses on the latter, we recognize that end-to-end planning benchmarks can also support research on LLM-generated models for external planners.
@@
LLMs-as-Planner Benchmarks: To determine whether testing PDDL domains have been leaked to training data of LLMs, Mystery Blocksworld (Kambhampati et al., 2024) obfuscates the classic Blocksworld (Gupta et al., 2010) planning problem by altering the named types so they are semantically equivalent but syntactically nonsensical. ALFWorld (Shridhar et al., 2021) and Household Guan et al. (2023) tackles the complexities of real-world typical household environment that uses PDDL semantics to produce textual observations and high-level actions. TravelPlanner (Xie et al., 2024a) assesses language models’ abilities in planning through agent-based interactions in a travel-planning environment. Zheng et al. (2024) extend this work with Natural Plan, which evaluates LLMs on realistic planning and scheduling benchmarks using APIs. PlanBench (Valmeekam et al., 2023a) aims to systematically evaluate LLM planning capabilities with an emphasis on cost-optimal planning and plan verification. ACPBench (Kokel et al., 2024) standardizes evaluation tasks and metrics for assessing reasoning about actions, changes (transitions), and planning—across 13 domains, on 22 SOTA language models. AutoPlanBench (Stein and Koller, 2023) first converts PDDL planning benchmarks into NL via LLMs, and then tasks LLMs to produce a plan through various prompting techniques. Bohnet et al. (2024) introduce a scalable benchmark suite in both PDDL and natural language to evaluate LLMs across diverse planning strategies, along with a method for translating PDDL benchmarks into natural language. Puerta-Merino et al. (2025) propose a road map and benchmark to address the gap of LLM integration in Hierarchical Planning (HP).
@@
LLMs-as-Planning-Formalizers Benchmarks: Planetarium (Zuo et al., 2024) provides a rigorous benchmark for evaluating PDDL task/problems produced by LLMs, highlighting two key issues: (i) LLMs can produce valid code that misaligns with the original NL description, and (ii) evaluation sets often use NL descriptions too similar to the ground truth, reducing the task’s challenge. The benchmark assesses LLMs’ ability to generate PDDL problems across varying levels of abstraction and size. However, it currently only supports Blocksworld, Gripper, and Floor Tile domains—well-known but narrow in dataset variability. On the other hand, Text2World (Hu et al., 2025) introduces an automated pipeline for domain extraction and rigorous multi-criteria metrics that address the limitations of narrow domain scope (Liu et al., 2023a) and indirect evaluation methods in end-to-end plan assessments. Key metrics, including executability, structural similarity, and component-wise F1 scores, are employed while exploring state-of-the-art LLMs and fine-tuning techniques. However, the reliance on executability as a gating metric excludes non-executable domains from component-wise scoring, causing minor syntax errors to skew overall quality assessments.
@@
Summary and Future Directions: Assessing the quality of LLM-generated PDDL models (Zuo et al., 2024; Hu et al., 2025; Oswald et al., 2024) has made significant progress toward rigorous evaluation; however, the rapid leakage of training data to LLMs remains a major challenge, with (Hu et al., 2025) reporting high contamination rates in the evaluation domains from (Guan et al., 2023). Future work should explore solutions for establishing dynamic benchmark standards for domains, actively involving the planning community in its ongoing refinement. Khandelwal et al. (2024) proposes a tool for generating diverse and complex planning domains, which could serve as a foundation for such a benchmark.
4 Language-to-Plan (L2P)